Selected financial data for Quick Sell, Inc., a retail store, appear as follows.

Year 2 Year 1
Sales (all on account) $ 781,000 $ 601,000
Cost of goods sold 489,000 431,000
Average inventory during the year 118,000 108,000
Average receivables during the year 150,000 100,000

  

a-1. Compute the gross profit percentage for both years. (Round your percentage answers to the nearest whole number. i.e. 0.1234 as 12%.)

a-2. Compute the inventory turnover for both years. (Round your answers to 1 decimal place.)

a-3. Compute the accounts receivable turnover for both years. (Round your answers to 1 decimal place.)

b. Which of the following show a positive or negative trend?
Answer #1

a-1. Gross profit percentage = Gross profit / Sales

Year 2 = (781,000-489000)/781,000 = 37%

Year 1 = (601,000-431000)/601,000 = 28%

a-2. Inventory turnover = Cost of goods sold / Average inventory during the year

Year 2 = 489000/118000 = 4.1

Year 1 = 431000/108000 = 4.0

a-3. Accounts receivable turnover = Sales (on account) / Average receivables during the year

Year 2 = 781000/150000 = 5.2

Year 1 = 601000/100000 = 6.0

b. Year 2 show a positive trend.